OpenStudy (anonymous):

what is the size of the population at t=1000days when the population size at t days is given by the function s(t)=Ae^(kt) s(0)=1000 s(500)=7000

OpenStudy (rogue):

The s(0) tells you the initial population, which is A.\[s(t)=1000e^{kt}\]Plug in t = 500 and s(500) = 7000 and solve for k.\[7000 = 1000e^{500k}\]\[k = \frac {1}{500} \ln 7 \approx 0.0038918203\]So your function is \[s(t) = 1000e^{0.0038918203t}\]You want to find the population when t = 1000 days, so plug that into your function.\[s(1000) = 1000e^{0.0038918203\times 1000} = 49000\]

OpenStudy (anonymous):

Thank u!

OpenStudy (anonymous):

should it not be ln 7000

OpenStudy (anonymous):

ok i got it! never mind

OpenStudy (rogue):

Your welcome :) There is an easier way to do this though :P from t = 0 to t = 500, the population grew 7 times. So the population grows 7 times every 500 days. From t= 500 to t = 1000, is another 500 days, so its going to be 7 times the population at t = 500, which would be 7000 x 7 = 49000. Much easier than doing all that exponential function stuff.

OpenStudy (anonymous):

i never looked at it that way! thank you!